OpenStudy (anonymous):

When f(x) = x4 + x2 – 4x + 6 is divided by ( x + 1) , the remainder is (A) 10 (B) 0 (C) 12 (D) 8

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Can you explain ?

OpenStudy (mathmate):

Substitute x=-1 into the polynomial, i.e. evaluate f(-1) and that would be the remainder.

OpenStudy (mathmate):

x=-1 comes from the factor (x+1)=0, so f(-1) = (-1)^4 + (-1)^2 – 4(-1) + 6 = 12

OpenStudy (anonymous):

whatever \[f(-1)\] is

OpenStudy (anonymous):

Why (x+1) should be 0 ?

OpenStudy (turingtest):

when the polynomial is divided by (x+1) we get a remainder of zero, which means the polynomial can be written in the form\[P(x)(x+1)=f(x)\]so if you wanted the zero's of f(x), x=-1 would be one of them, but there is nothing in the problem to suggest that since the polynomial is a function and not equal to zero.

OpenStudy (turingtest):

P(x) is the polynomial that you get after dividing by (x+1)

OpenStudy (anonymous):

\[f(x) = x4 + x2 – 4x + 6=(x+1)\times (\text{ something}) + \text{remainder}\] \[f(-1)=(-1+1)\times (\text{ something}) + \text{remainder}\] \[f(-1)=(0)\times (\text{ something}) + \text{remainder}\] \[f(-1)=\text{ remainder}\]

OpenStudy (turingtest):

can you do synthetic division @abdul shabeer?

OpenStudy (anonymous):

What does that mean ?

OpenStudy (mr.math):

Nice @satellite! :)

OpenStudy (turingtest):

it's a fast way of dividing polynomials by expressions of the form (x-a) if I divide some polynomial P(x) by (x-a) and get no remainder, then a is a root of the equation. http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

Yes, I know this division

OpenStudy (anonymous):

12